OpenStudy (anonymous):

What is the slope of the line given by the equation below? y + 2 = -3(x - 5) A. 3 B. 2 C. -5 D. -3

OpenStudy (haseeb96):

d is the correct answer

OpenStudy (anonymous):

Thanks if you mind me asking how did you get that answer? @Haseeb96

OpenStudy (haseeb96):

ax+by+c =0 if this type of equation is given to u then use this formula m= - a/b and if this type by = ax-c then the term along x is yur slope

OpenStudy (anonymous):

Alright , Thank you so much :) . @Haseeb96

OpenStudy (haseeb96):

:) just pray for me to get admission in university then i will teach math in yu difficult areas according to ur question

OpenStudy (anonymous):

Of course :) Lots of luck @Haseeb96